OpenStudy (anonymous):

Simplify the expression. 57 – 12 ÷ 4 • 3

OpenStudy (mathstudent55):

Do you know the order of operations?

OpenStudy (anonymous):

Order of Operations Multiplication and Division, then addition and subtraction. 57 - 12/4 *3 57 - 12/12 57 - 1 56

OpenStudy (anonymous):

nvm i got it :P

OpenStudy (anonymous):

56 is wrong lol, its 48

OpenStudy (mathstudent55):

57 – 12 ÷ 4 • 3 = 57 – 3 • 3 = 57 – 9 = 48